# Introducing NewCluster: the first half of the history of a high-resolution cluster simulation

### v1

@@ 2.3.1 Stellar Physics and Chemical Evolution
The star formation rate (SFR) in the simulation is calculated based on the gravo-thermo-turbulent model of Kimm et al. (2017) without applying the turbulent Jeans length criterion, thereby preserving scale invariance, as in Dubois et al. (2021). The exact formation efficiency follows the equation provided by Federrath & Klessen (2012). In each time step, the amount of stellar mass formed is checked within each gas cell above a star formation density threshold of $n_{\star}=5\,\rm H\,cm^{-3}$ to form a star particle.
@@
The stellar feedback in the simulation consists of two kinds: supernovae (SNe) and stellar winds. The SN feedback is implemented as two independent types, Type II and Type Ia (as SNII and SNIa, respectively), both of which are implemented as the form of mechanical feedback scheme of Kimm & Cen (2014) that takes account of the empirical boost depending on the local density of the explosion (Kimm et al., 2017). Every star particle that has reached the age of 5 Myr triggers the SNII explosions. The produced feedback energy per mass is set to be $e_{\rm SNII}=0.045\times 10^{51}\,{\rm erg}\,M_{\sun}^{-1}$.
The frequency of SNIa is set to follow the delay time distribution (DTD) in the form of a power law (Maoz et al., 2012),
@@
For a realistic production of chemical elements within galaxies, we have included mass return from stellar ejecta of SNe type II and Ia, and stellar winds in the simulation. Throughout the simulation, we assume a Chabrier (2003) initial mass function for the integration of chemical yield from the stellar population. Considering each stellar particle as a simple stellar population (SSP), chemical enrichment for ten species (H, D, He, C, N, O, Mg, Fe, Si, and S) is computed using the Starburst99 code (Leitherer et al., 1999, 2014). [^2] D denotes deuterium that cannot be produced from ejecta of stellar origin, but only be present as an initial fraction of $({\rm D}/{\rm H})=2.6\times 10^{-5}$ (Cyburt et al., 2016). Helium is not explicitly traced in the simulation, but is considered as a residual component after the nine tracked elements, with an initial abundance of $Y=0.2477$ (Peimbert et al., 2007). For the other heavier elements, the gas is set to be metal-free at the initial state. Yields from stellar winds are derived assuming the Geneva stellar wind model (Schaller et al., 1992; Maeder & Meynet, 2000). At the same time, energy produced from stellar wind is directly injected into the grid where the star is positioned. SNII explosions also inject newly synthesised elements into surroundings based on the tabulated value for the SNII yields in Kobayashi et al. (2006), assuming that each SSP has the progenitor mass range of $8-50\,M_{\odot}$, which corresponds to the ejecta mass fraction for each stellar particle is $M_{\rm ejecta}/M_{*}=0.311$. The yield from SNIa is derived from Iwamoto et al. (1999) with the mass of the ejecta in every burst corresponding to the Chandrasekhar limit ($\sim 1.4\,M_{\sun}$), with corresponding decline of mass in each stellar particle.
@@ 2.3.2 Dust Formation and Evolution
The simulation also computes dust formation and evolution as a passive scalar that evolves on-the-fly throughout cosmic time. The dust model implemented for NewCluster is based on that described in Dubois et al. (2024) (which we refer to for further technical details) and accounts for several key physical processes, including dust formation, growth, destruction, and size evolution. Here, we provide a brief summary of the methodology.
Dust grain species are modeled with two distinct characteristic chemical compositions: carbonaceous and silicate grains. We assume that silicate grains consist of a single form of amorphous olivine, with compound MgFeSiO4, and that carbonaceous grains are composed purely of carbon. To follow the grain size distribution of dust, the two different chemical compositions are further decomposed into two different grain sizes (adopting the two-size bin approximation of Hirashita, 2015) of radius 5 nm for small grains and 0.1 $\mu$m for large grains.
@@
For dust grain production, we assume two channels: the formation from the stellar ejecta and growth through the accretion of gas-phase metals in the ISM. We consider the dust production from SNe type II and Ia, and asymptotic giant branch (AGB) winds based on the feedback models and stellar yields described above with different dust condensation efficiencies that depend on the stellar origin and on the C/O ratio (see Dubois et al., 2024 for the exact values). Accretion of gas-phase metals onto dust grains in dense and cold ISM regions is also implemented as a subgrid model. Dust accretion rates are estimated based on the gas density, temperature, and metallicity in each gas cell and depend on the grain size and its material density ($s_{\rm gr}=3.3$ and $2.2\,\rm g\,cm^{-3}$ for silicate and carbonaceous grains respectively). The local gas velocity dispersion is used to sample a subgrid log-normal shape of the probability density function of the gas density, assuming isothermal turbulence. We assume that the gas-phase accretion occurs only for (subgrid) gas densities lower than $n_{\rm thr}=10^{4}\,{\rm H\,cm^{-3}}$, considering the formation of ice mantles on grain surfaces, which inhibit metal accretion onto pre-existing grain cores (Cuppen & Herbst, 2007; Hollenbach et al., 2009). The Coulomb enhancement factor, which accounts for the electrostatic effects acting differently on various grain species, is not included in the NewCluster dust model. For dust destruction, we include thermal sputtering and supernova-driven destruction from both Ia and II explosions with increased efficiencies with decreasing grain sizes. The thermal sputtering timescale with temperature is estimated following Tsai & Mathews (1995). SNe destroy the enclosed dust until the shock velocity decreases below the activation threshold of $100\,\rm km\,s^{-1}$, with identical destruction efficiencies for graphite and silicate grains. Astration completely destroys the dust engulfed in the star formation process.
@@
The two grain sizes of each grain chemical composition can exchange mass by shattering and coagulation, respectively, decreasing and increasing the grain sizes. Shattering operates at intermediate gas densities ($n\lesssim 10\,\rm H\,cm^{-3}$) where grain turbulent velocities ($\sigma_{\rm gr}\gtrsim 10\,\rm km\,s^{-1}$) are sufficiently high to fragment the grains (Jones et al., 1996; Yan et al., 2004). In contrast, coagulation takes place in the dense gas phase ($n\gtrsim 100\,\rm H\,cm^{-3}$) where turbulence is sufficiently low ($\sigma_{\rm gr}\lesssim 1\,\rm km\,s^{-1}$) to preserve the grain structure and to allow grain size growth by collisions (Chokshi et al., 1993; Poppe & Blum, 1997). The shattering and coagulation timescales are calculated following Granato et al. (2021) and Aoyama et al. (2017), respectively. Readers are referred to Dubois et al. (2024) for further details.
@@ 2.3.3 Massive Black Hole Physics
For AGN feedback, we employ the dual mode feedback scheme including the spin-evolution of massive black holes (MBHs) (Dubois et al., 2012, 2014b, 2021). The formation criterion of seed MBHs is applied based on the gas and stellar properties in each cell. The target cell should have a gas density higher than $n_{\rm BH}=100\,\rm H\,cm^{-3}$, and a stellar mass within the cell higher than $1.88\times 10^{7}M_{\odot}$. This corresponds to a density of $2500\,\rm H\,cm^{-3}$ at the best spatial resolution at $z=0$. The motivation for using the stellar mass of the cell instead of its density is to make the occurrence frequency of peak-stellar-density cells independent of spatial resolution. Since NewCluster employs a time-dependent minimum spatial resolution, this approach helps the formation of seed black holes to be more uniformly distributed throughout the epochs. We also require the stellar velocity dispersion to be $\sigma_{*}>50\,\rm km\,s^{-1}$ and the stellar age to be $t_{*}>50\,{\rm Myr}$ to avoid selecting short-lived young clumps as seed MBH formation sites.
@@
The exclusion radius, which is the radius of the spherical region that prevents the formation of new MBH around the pre-existing MBH, is set to be $4\,\,\mathrm{kpc}$. This constraint was originally imposed to prevent the overproduction of seeds inside a single galaxy when the formation criterion was easily met. Compared to previous RAMSES simulations, our threshold scheme ensures the formation site of seeds to be reliably dense stellar concentrations that last long enough, making a large exclusion radius (e.g., 50 comoving kpc in Dubois et al., 2021) unnecessary. Our choice of $4\,kpc$ is enough to prevent the formation of multiple seeds near the galactic centre.
@@
Dynamical friction of MBHs (Chandrasekhar, 1943; Ostriker, 1999) is implemented for gas from Dubois et al. (2014b) and stars from Pfister et al. (2019). We find that seed-mass MBHs do not settle with their own dynamical friction, even when they are near the central region of the galaxy, resulting in permanent oscillation near the galaxy centre. This issue has also been reported in other simulations (Ma et al., 2021; Bahé et al., 2022). To alleviate the problem, artificial acceleration and repositioning of the MBH have been utilised, for example, by boosting dynamical friction or the accretion rate of the MBH by a factor that depends on the gas density (Weinberger et al., 2017; Wellons et al., 2023; Tremmel et al., 2018; Dubois et al., 2021).
@@
MBHs seem to show stable settling even in higher-resolution simulations that take into account dynamical interactions with both the stars and the gas components within the Bondi-Hoyle radius (Pfister et al., 2019; Lescaudron et al., 2023; Li et al., 2024). Other simulations address stellar systems associated with the MBH (e.g., nuclear star clusters) as a critical player for the settling of the MBH (Biernacki et al., 2017; Ogiya et al., 2020; Mukherjee et al., 2025; Zhou et al., 2025; Shin et al., 2025). Since our 686868\,68pc spatial resolution is not enough to capture these behaviours, we apply a mild boost factor on the stellar dynamical friction of the MBH:
@@ 2.3.4 Heating and Cooling of Gas
The general radiative heating and cooling model follows the framework of Dubois et al. (2021). The simulation employs a UV background from Haardt & Madau (1996) with reionisation redshift $z_{\rm reion}=10$. Gas with density $n>0.01\,{\rm H\,cm^{-3}}$ (Rosdahl & Blaizot, 2012) has a reduced UV photo-heating rate due to self-shielding by a factor of $\exp[-n_{\rm H}/(0.01\,{\rm H\,cm^{-3}})]$. We used tabulated cooling rates from Sutherland & Dopita (1993) and Dalgarno & McCray (1972) that vary with the metallicity which is defined as the sum of the abundances of all gas-phase elements excluding H, He, and D.
The simulation includes an additional cooling effect from collisional interactions between electrons and dust grains at high temperature. We adopt the cooling rates from Dwek & Werner (1981), which depend on the gas temperature, dust grain size, and the number densities of hydrogen and electrons, assuming that the gas is fully ionised. The dust contribution to the gas cooling rate significantly decreases below $\sim 10^{6}\,\rm K$, thereby preventing dust from affecting the contraction of cold gas at lower temperatures (see Dubois et al., 2024 for details).
@@ 2.4 Halo and Galaxy Detection
Galaxies and halos are identified separately using the AdaptaHOP algorithm (Aubert et al., 2004; Tweed et al., 2009). This method detects peaks and saddle points in the particle density field, which are used to estimate the initial centres and boundaries of halos. To build a substructure hierarchy, we adopt the “most massive sub-node method,” which regards the most massive (leaf) substructures as the central part of the host halo while other substructures inside the host halo are classified as “subhalos” (see Tweed et al., 2009 for details). The minimum number of particles is set to 100, corresponding to $\sim 10^{6}\,M_{\odot}$ and $\sim 10^{8}\,M_{\odot}$ for galaxies and halos, respectively. In total, 5,570 galaxies and 89,994 halos are identified at $z=0.8$. However, as the zoom-in simulation contains DM particles with different resolutions, some halos are contaminated by low-resolution particles. To consider this, we applied a strict condition that excludes halos with at least one poor-resolution DM particle, resulting in 73,745 halos in total. The mass ranges are $10^{6.26\text{--}12.12}\,M_{\odot}$ for galaxies and $10^{8.12\text{--}14.15}\,M_{\odot}$ for halos. The two most massive halos, which are merging at $z=0.8$, also satisfy this condition and contain no low-resolution DM particles.
@@
Although AdaptaHOP computes several main halo properties such as virial radius [^3] In AdaptaHOP, it directly computes the virial condition using member particles and defines the virial radius ($R_{\rm vir}$) where inside particles physically satisfy the condition. or spin parameter, we further calculated key quantities related to stellar or gas components. To measure these quantities, we apply different strategies for stellar and gas properties. For stellar properties of a galaxy (e.g., SFR or metallicity), we only consider member particles identified by the galaxy finder. On the other hand, gas properties (e.g., gas mass) are measured using several radius cuts, such as the stellar half-mass radius $R_{50}$ (or $R_{90}$), the radius within which a galaxy contains 50% (90%) of its stellar mass, and $R_{\rm max}$, the distance from the centre to the farthest member particle. We publicly provide all value-added catalogues on the simulation website. [^4] https://gemsimulation.com

### v2

@@ 2.3.1 Stellar physics and chemical evolution
The star formation rate (SFR) in the simulation is calculated based on the gravo-thermo-turbulent model of 2017MNRAS.466.4826K without applying the turbulent Jeans length criterion, thereby preserving scale invariance, as in 2021A&A...651A.109D. The exact formation efficiency follows the equation provided by 2012ApJ...761..156F (The multi-ff PN model). In each time step, the amount of stellar mass formed is checked within each gas cell above a star formation density threshold of $n_{\star}=5\,\rm H\,cm^{-3}$ to form a star particle.
@@
The stellar feedback in the simulation consists of two kinds: supernovae (SNe) and stellar winds. The SN feedback is implemented as two independent types, Type II and Type Ia (as SNII and SNIa, respectively), both of which are implemented as the form of mechanical feedback scheme of 2014ApJ...788..121K that takes account of the empirical boost depending on the local density of the explosion (2017MNRAS.466.4826K). Every star particle that has reached the age of 5 Myr triggers the SNII explosions. The produced feedback energy per mass is set to be $e_{\rm SNII}=0.045\times 10^{51}\,{\rm erg}\,M_{\sun}^{-1}$.
The frequency of SNIa is set to follow the delay time distribution (DTD) in the form of a power law (Maoz12),
@@
For a realistic production of chemical elements within galaxies, we have included mass return from stellar ejecta of SNe type II and Ia, and stellar winds in the simulation. Throughout the simulation, we assume a 2003PASP..115..763C initial mass function for the integration of chemical yield from the stellar population. Considering each stellar particle as a simple stellar population (SSP), chemical enrichment for ten species (H, D, He, C, N, O, Mg, Fe, Si, and S) is computed using the Starburst99 code (1999ApJS..123....3L; 2014ApJS..212...14L). [^2] D denotes deuterium that cannot be produced from ejecta of stellar origin, but only be present as an initial fraction of $({\rm D}/{\rm H})=2.6\times 10^{-5}$ (2016RvMP...88a5004C). Helium is not explicitly traced in the simulation, but is considered as a residual component after the nine tracked elements, with an initial abundance of $Y=0.2477$ (2007ApJ...666..636P). For the other heavier elements, the gas is set to be metal-free at the initial state. Yields from stellar winds are derived assuming the Geneva stellar wind model (Schaller92; Maeder00). At the same time, energy produced from stellar wind is directly injected into the grid where the star is positioned. SNII explosions also inject newly synthesised elements into surroundings based on the tabulated value for the SNII yields in 2006ApJ...653.1145K, assuming that each SSP has the progenitor mass range of $8-50\,M_{\odot}$, which corresponds to the ejecta mass fraction for each stellar particle is $M_{\rm ejecta}/M_{*}=0.311$. The yield from SNIa is derived from 1999ApJS..125..439I with the mass of the ejecta in every burst corresponding to the Chandrasekhar limit ($\sim 1.4\,M_{\sun}$), with corresponding decline of mass in each stellar particle.
@@ 2.3.2 Dust formation and evolution
The simulation also computes dust formation and evolution that evolves on-the-fly throughout cosmic time. The dust model implemented for NewCluster is based on that described in Yohan24 (which we refer to for further technical details) and accounts for several key physical processes, including dust formation, growth, destruction, and size evolution. Here, we provide a brief summary of the methodology.
[…]
@@
For dust grain production, we assume two channels: the formation from the stellar ejecta and growth through the accretion of gas-phase metals in the ISM. We consider the dust production from SNe type II and Ia, and asymptotic giant branch (AGB) winds based on the feedback models and stellar yields described above with different dust condensation efficiencies that depend on the stellar origin and on the C/O ratio (see Yohan24 for the exact values). Accretion of gas-phase metals onto dust grains in dense and cold ISM regions is also implemented as a subgrid model. Dust accretion rates are estimated based on the gas density, temperature, and metallicity in each gas cell and depend on the grain size and its material density ($s_{\rm gr}=3.3$ and $2.2\,\rm g\,cm^{-3}$ for silicate and carbonaceous grains respectively). The local gas velocity dispersion is used to sample a subgrid log-normal shape of the probability density function of the gas density, assuming isothermal turbulence. We assume that the gas-phase accretion occurs only for (subgrid) gas densities lower than $n_{\rm thr}=10^{4}\,{\rm H\,cm^{-3}}$, considering the formation of ice mantles on grain surfaces, which inhibit metal accretion onto pre-existing grain cores (Cuppen&Herbst07; Hollenbach09). The Coulomb enhancement factor, which accounts for the electrostatic effects acting differently on various grain species, is not included in the NewCluster dust model. For dust destruction, we include thermal sputtering and supernova-driven destruction from both Ia and II explosions with increased efficiencies with decreasing grain sizes. The thermal sputtering timescale with temperature is estimated following Tsai&Mathews95. SNe destroy the enclosed dust until the shock velocity decreases below the activation threshold of $100\,\rm km\,s^{-1}$, with identical destruction efficiencies for graphite and silicate grains. Astration completely destroys the dust engulfed in the star formation process.
@@
The two grain sizes of each grain chemical composition can exchange mass by shattering and coagulation, respectively, decreasing and increasing the grain sizes. Shattering operates at intermediate gas densities ($n\lesssim 10\,\rm H\,cm^{-3}$) where grain turbulent velocities ($\sigma_{\rm gr}\gtrsim 10\,\rm km\,s^{-1}$) are sufficiently high to fragment the grains (Jones1996; Yan04). In contrast, coagulation takes place in the dense gas phase ($n\gtrsim 100\,\rm H\,cm^{-3}$) where turbulence is sufficiently low ($\sigma_{\rm gr}\lesssim 1\,\rm km\,s^{-1}$) to preserve the grain structure and to allow grain size growth by collisions (Chokshi1993; Poppe&Blum1997). The shattering and coagulation timescales are calculated following Granato21 and Aoyama17, respectively. Throughout the simulation, the dust components are advected with the gas flow, behaving as passive scalars. Readers are referred to Yohan24 for further details.
@@ 2.3.3 Massive black hole physics
For AGN feedback, we employ the dual mode feedback scheme including the spin-evolution of massive black holes (MBHs) (2012MNRAS.420.2662D; 2014MNRAS.440.1590D; 2021A&A...651A.109D). We set the thermal feedback efficiency for the AGN to 0.05, which is three times lower than in 2021A&A...651A.109D. This reduces the amount of feedback energy released per unit accreted mass in quasar mode, thereby promoting black hole growth and increasing their mass at a given stellar mass, which was underestimated in previous simulations relative to observations. The formation criterion of seed MBHs is applied based on the gas and stellar properties in each cell. The target cell should have a gas density higher than $n_{\rm BH}=100\,\rm H\,cm^{-3}$, and a stellar mass within the cell higher than $1.88\times 10^{7}M_{\odot}$. This corresponds to a density of $2500\,\rm H\,cm^{-3}$ at the best spatial resolution at $z=0$. The motivation for using the stellar mass of the cell instead of its density is to make the occurrence frequency of peak-stellar-density cells independent of spatial resolution. Since NewCluster employs a time-dependent minimum spatial resolution, this approach helps the formation of seed black holes to be more uniformly distributed throughout the epochs. We also require the stellar velocity dispersion to be $\sigma_{*}>50\,\rm km\,s^{-1}$ and the stellar age to be $t_{*}>50\,{\rm Myr}$ to avoid selecting short-lived young clumps as seed MBH formation sites.
@@
The exclusion radius, which is the radius of the spherical region that prevents the formation of new MBH around the pre-existing MBH, is set to be $4\,\,\mathrm{kpc}$. This constraint was originally imposed to prevent the overproduction of seeds inside a single galaxy when the formation criterion was easily met. Compared to previous RAMSES simulations, our threshold scheme ensures the formation site of seeds to be reliably dense stellar concentrations that last long enough, making a large exclusion radius (e.g., 50 comoving kpc in 2021A&A...651A.109D) unnecessary. Our choice of 4kpc4\,kpc is enough to prevent the formation of multiple seeds near the galactic centre.
@@
Dynamical friction of MBHs (1943ApJ....97..255C; 1999ApJ...513..252O) is implemented for gas from 2014MNRAS.440.1590D and stars from 2019MNRAS.486..101P. We find that seed-mass MBHs do not settle with their own dynamical friction, even when they are near the central region of the galaxy, resulting in permanent oscillation near the galaxy centre. This issue has also been reported in other simulations (Ma+2021; Bahe+2022). To alleviate the problem, artificial acceleration and repositioning of the MBH have been utilised, for example, by boosting dynamical friction or the accretion rate of the MBH by a factor that depends on the gas density (Weinberger+17; Wellons+23; Tremmel2018; 2021A&A...651A.109D).
@@
MBHs seem to show stable settling even in higher-resolution simulations that take into account dynamical interactions with both the stars and the gas components within the Bondi-Hoyle radius (2019MNRAS.486..101P; 2023A&A...674A.217L; 2025A&A...701A.232L). Other simulations address stellar systems associated with the MBH (e.g., nuclear star clusters) as a critical player for the settling of the MBH (2017MNRAS.469..295B; 2020MNRAS.493.3676O; 2025ApJ...981..203M; 2025ApJ...980...79Z; Shin25). Since our 6868\,pc spatial resolution is not enough to capture these behaviours, we apply a mild boost factor on the stellar dynamical friction of the MBH:
@@ 2.3.4 Heating and cooling of gas
The general radiative heating and cooling model follows the framework of 2021A&A...651A.109D. The simulation employs a UV background from 1996ApJ...461...20H with reionisation redshift $z_{\rm reion}=10$. Gas with density $n>0.01\,{\rm H\,cm^{-3}}$ (2012MNRAS.423..344R) has a reduced UV photo-heating rate due to self-shielding by a factor of $\exp[-n_{\rm H}/(0.01\,{\rm H\,cm^{-3}})]$. We used tabulated cooling rates from 1993ApJS...88..253S and 1972ARA&A..10..375D that vary with the metallicity which is defined as the sum of the abundances of all gas-phase elements excluding H, He, and D.
The simulation includes an additional cooling effect from collisional interactions between electrons and dust grains at high temperature. We adopt the cooling rates from 1981ApJ...248..138D, which depend on the gas temperature, dust grain size, and the number densities of hydrogen and electrons, assuming that the gas is fully ionised. The dust contribution to the gas cooling rate significantly decreases below $\sim 10^{6}\,\rm K$, thereby preventing dust from affecting the contraction of cold gas at lower temperatures (see Yohan24 for details).
@@ 2.4 Halo and galaxy detection
Galaxies and halos are identified separately using the AdaptaHOP algorithm (Aubert2004; Tweed2009). This method detects peaks and saddle points in the particle density field, which are used to estimate the initial centres and boundaries of halos. To build a substructure hierarchy, we adopt the “most massive sub-node method,” which regards the most massive (leaf) substructures as the central part of the host halo while other substructures inside the host halo are classified as “subhalos” (see Tweed2009 for details). The minimum number of particles is set to 100, corresponding to $\sim 10^{6}\,M_{\odot}$ and $\sim 10^{8}\,M_{\odot}$ for galaxies and halos, respectively. In total, 5,570 galaxies and 89,994 halos are identified at $z=0.8$. However, as the zoom-in simulation contains DM particles with different resolutions, some halos are contaminated by low-resolution particles. To consider this, we applied a strict condition that excludes halos with at least one poor-resolution DM particle, resulting in 73,745 halos in total. The mass ranges are $10^{6.26\text{--}12.12}\,M_{\odot}$ for galaxies and $10^{8.12\text{--}14.15}\,M_{\odot}$ for halos. The two most massive halos, which are merging at $z=0.8$, also satisfy this condition and contain no low-resolution DM particles.
@@
Although AdaptaHOP computes several main halo properties such as virial radius [^3] In AdaptaHOP, it directly computes the virial condition using member particles and defines the virial radius ($R_{\rm vir}$) where inside particles physically satisfy the condition. or spin parameter, we further calculated key quantities related to stellar or gas components. To measure these quantities, we apply different strategies for stellar and gas properties. For stellar properties of a galaxy (e.g., SFR or metallicity), we only consider member particles identified by the galaxy finder. On the other hand, gas properties (e.g., gas mass) are measured using several radius cuts, such as the stellar half-mass radius $R_{\rm 50,3D}$ (or $R_{\rm 90,3D}$), the 3-dimensional radius within which a galaxy contains 50% (90%) of its stellar mass, and $R_{\rm max}$, the distance from the centre to the farthest member particle. We publicly provide all value-added catalogues on the simulation website. [^4] https://gemsimulation.com